OpenStudy (anonymous):

what is 811 divided 2 estimating

TheSmartOne (thesmartone):

Round 811 to 810 810/2=?

TheSmartOne (thesmartone):

So 810/2=405

OpenStudy (anonymous):

If you give a very rough estimation, then you'd say 811 ~ 800, and you get 811/2 ~ 800/2 = 400

TheSmartOne (thesmartone):

true^^